### v1

@@ 5.6 Ablation Studies
Keypoints. Not all joints of the hands are equally informative for understanding hand actions. For instance, fingertips exhibit greater mobility compared to the inner joints. Moreover, from an egocentric viewpoint, certain joints are more prone to self-occlusion than others. In Tab. 5, we present the impacts of incorporating varying numbers of joints on verb recognition within the Assembly101 dataset [51]. For the case of 6 joints, we consider only the wrist joint along with the five fingertips, and to expand to 11 joints, we additionally incorporate all the joints along the index finger and thumb, which are least affected by self-occlusion. We also show the effect of including Global Wrist Token, which acts as a reference to the global motion of the hands while encoding micro-actions.
@@
Method Action Verb Object TSM egocentric (fuse 4 views) 33.80 59.00 46.50 Egocentric (e3) + Pose 36.07 65.52 45.82 Egocentric (e4) + Pose 35.56 65.79 45.20 Egocentric (e3$+$e4) + Pose 38.05 66.32 47.86 Table 4: Egocentric action recognition in Assembly101 [51]. TSM features from [51] are used as RGB frame features.  #Joints Global Verb  Reference Accuracy (%) 21 64.17 21 64.90 11 64.77 6 63.70 Table 5: Keypoint ablation for verb recognition in Assembly101 [51].
#Frames 1 15 30 60 120 Verb  59.12  63.70  63.68  63.51  62.29 Accuracy(%) Table 6: Micro-action length ablation for verb recognition in Assembly101 [51]. Temp. Agg. TCN LSTM Transformer Verb  62.95 63.34 63.70 Accuracy(%) Table 7: Ablating unimodal temporal aggregation choices for verb recognition in Assembly101 [51].
@@
Micro-action length. As the resized input comprises a set number of pose frames, enlarging the window size for a micro-action will lead to a decrease in the number of micro-actions to aggregate, and conversely. In Tab. 6, we vary the micro-action length for verb recognition in the Assembly101 dataset [51] using 6 joints per hand as input, i.e. fingertips and wrist joint. The input pose clip is temporally resized to $T^{\prime}=120$ before breaking into micro-actions. Lengths 1 and 120 represent two extreme versions with frame-based and trajectory-based encoding, respectively, while the others conform to our micro-action-based formulation.
@@
Temporal Aggregation. After extracting features for micro-actions, aggregation for action recognition can be done with any sequence model. In Tab. 7, we evaluate the effectiveness of different temporal aggregation methods for verb recognition in the Assembly101 dataset [51]. Here, we use 6 joints per hand as input, i.e. fingertips and wrist joint.
@@
Loss components. To assess the individual contributions of different components, we begin with a basic configuration. We then systematically introduce each element to understand its impact on the overall performance as observed in Tab. 8. Incorporating modality interaction between RGB and pose at the micro-action level through a multimodal tokenizer enhances action accuracy. The introduction of auxiliary losses also has a positive impact, resulting in an overall improvement of 2.08% for Assembly101 [51] and 7.45% for H2O [33].
@@
 Multimodal Reconstruction  Verb & Object Action Accuracy (%) Tokenizer Loss Loss Assembly101 H2O 38.98 85.95 40.19 88.84 40.24 89.26 40.56 90.50 41.06 93.39
Table 8: Ablating tokenization and different losses for action recognition on Assembly101 [51] and H2O [33].
@@ 6 Conclusion
With the growing interest in AR/VR and wearables, hand pose estimation has rapidly advanced, leading to the development of dedicated hardware independent of vision. As such, hand poses hold promise as a compact, domain-independent, rich modality to complement computer vision in the near future. To address the under-explored domain of using 3D hand poses as a modality for hand-object interaction recognition, we highlight the fundamental differences between hand poses and full-body human skeletons. We introduce HandFormer, a novel multimodal transformer that leverages dense sequences of 3D hand poses with sparsely sampled RGB frames to achieve state-of-the-art action recognition performance. Our model also reduces computational requirements, offering immediate significance across various low-resource applications in mobile devices.
Limitations. Our method relies on the availability of hand poses, which, if extracted from the visual modality with pose estimation tools [26, 27], can encounter out-of-view scenarios and produce noisy poses. Although our experiments show that such estimated poses can still achieve good accuracy, further research can be done to explicitly address this phenomenon. We also assume that uniform sampling of RGB frames from each micro-action should provide us with good representations for understanding the semantic context, which is not true as all frames are not equally important in action. In such a case, adaptive frame sampling methods can be employed which we leave for future work.
@@
A Statistical Analysis: Full-body Skeleton vs. Hand
@@
This section serves as an extension of Sec. 3 to statistically analyze the difference between hand poses and full-body poses. Action recognition datasets [53, 40], which primarily focus on full-body actions, often include actions involving partial body movements. These actions exhibit limited motion when viewed with respect to the entire skeleton, resulting in one or more relatively static joints. The change in the locations of moving joints with respect to such static joints can provide useful action cues. However, hand motions typically feature no such static reference points, as all the hand joints move together most of the time, making small changes in articulation to perform an action. To illustrate this difference, we randomly sample 1000 pose sequences from NTU RGB+D 120 [40] (full-body) and Assembly101 [51] (hands). We take the distance covered between two consecutive frames for each joint to form a distance array for the corresponding joint $j$ in a given sequence, which is determined by:
@@
[…]
For all the sampled 1000 sequences, we take the temporal sequences $\{d_{j_{sta}}(t)\}_{t=1}^{T}$ and $\{d_{j_{dyn}}(t)\}_{t=1}^{T}$, normalize the distance values using the diameter of the corresponding skeleton, and plot the sequences separately in Fig. 7. As can be observed, compared to the distances covered by the most active joints in NTU RGB+D 120, the least active joints show significantly lower movement, effectively serving as the static reference points. On the other hand, the distinction in distance arrays between the most and the least active joints in Assembly101 is less pronounced. In addition, we calculated the Pearson correlation coefficient, denoted as $r$, between $\{d_{j_{sta}}(t)\}_{t=1}^{T}$ and $\{d_{j_{dyn}}(t)\}_{t=1}^{T}$ for all Assembly101 sequences, resulting in a value of 0.93. Conversely, for NTU RGB+D 120, the corresponding correlation coefficient is 0.33. This suggests strong coupling among hand joints during motion, and full-skeleton movement is more dominant in hand poses compared to full-body poses. Consequently, modeling dependencies between spatiotemporally distant joints is less effective for the highly dynamic hand motion (also discussed in Sec. 3). Therefore, by considering both long-term motion patterns and short-term articulation changes, our method facilitates efficient spatiotemporal factorization through micro-actions (Sec. 4.1). We also incorporate the full-skeletal motion from the entire action during micro-action encoding, using a global wrist token as a reference (Sec. 4.2).
@@
Fig. 8 illustrates sample heatmaps from NTU RGB+D 120 [40] and Assembly101 [51]. Keypoints in full-body human poses are often prominently situated, with minimal self-occlusion, and the subject is typically centered within the frame. As viewed in Fig. 7(a), reducing pose dimensions to 2D does not significantly compromise detail; rather, it enhances input reliability by simplifying the pose estimation problem. However, this advantage diminishes when applied to hand poses. Hand poses present unique challenges, such as frequent self-occlusion and closer proximity of the keypoints, which are exacerbated by reducing the dimension to 2D. To empirically analyze this phenomenon, we evaluate HandFormer-B/6 with 2D and 3D poses for recognizing verbs in Assembly101[51] and report in Tab. 9. This analysis reveals about $5$% difference in favor of the 3D input. Furthermore, PoseConv3D [20] introduces a CNN-based approach with 2D keypoints, which directly utilizes heatmaps from the pose estimator or generates Gaussian heatmaps from the 2D coordinates. However, feeding heatmaps to the model can diminish the clarity of keypoints, particularly when they are in close proximity, as is often the case with hand poses. Hence, PoseConv3D [20] performs poorly in recognizing hand actions, as evident in Tab. 9.
@@
C Efficient Alternative for Frame Encoder
Except for our experiments on Assembly101 [51] with monochrome egocentric videos (Sec. 5.5) , all our multimodal results are obtained using DINOv2 [46] features extracted from the input RGB frames. This approach enables us to assess the effectiveness of image-based foundation models in videos, leveraging all-purpose features from RGB frames, also achieving strong performance in cross-view generalization (Sec. 5.4). While using such frozen encoders for feature extraction can expedite training and mitigate the need for domain-specific modeling, the inference can be compute-heavy due to the large ViT backbones of foundation models. To alleviate this, we propose a pretraining scheme that allows us to use a ResNet50 [28] backbone to extract image features, thereby reducing inference time. Specifically, we first train a TSM [39] model with a ResNet50 [28] backbone for action recognition, utilizing all action clips and then dropping the classification layer. This ResNet50 becomes the frozen image encoder in our proposed architecture, replacing ViT. During the training and inference of HandFormer, this TSM backbone operates as a true image model (ResNet50), as we employ it on individual frames without any channel shifting. The TSM features provided in the Assembly101 [51] are generated in this way, and we utilize them in our egocentric action recognition experiments (Sec. 5.5).
@@ C Efficient Alternative for Frame Encoder
D Maintaining High Temporal Resolution at Low Cost
Our method is designed to perform action recognition efficiently in hand-object interaction videos. Obtaining efficiency in such a setup is challenging as we need to maintain a high temporal resolution to understand fine-grained hand movements that constitute the action. Therefore, we propose HandFormer using densely sampled pose frames and sparse RGB frames. In this section, we quantify the efficiency of this method compared to an alternative video model. As mentioned, understanding fine-grained hand motion demands a high temporal resolution to differentiate verb classes. For instance, relying on sparsely sampled frames may make actions like “screwing" and “unscrewing" indistinguishable. However, adopting a high temporal resolution with video models operating on RGB frames is challenging, primarily due to (i) the excessive computation associated with performing spatiotemporal operations on numerous frames, and (ii) the need to address redundancy in RGB frames to extract meaningful information.
@@ E.2 Hand-Object Interaction Crop
In hand-object interaction (HOI) videos, the region of interest typically centers around the hands, capturing crucial information about the interacting object and the type of interaction. Leveraging 3D hand poses obtained through a readily available pose estimator [26], we project these poses onto RGB frames, extract the enclosing rectangle of the projected 2D pose, and expand it by 25% to define the ROI crop. However, relying solely on the cropped region can occasionally mislead the model for three potential reasons: i) failure of the pose estimator on certain frames, leading to the absence of useful features from the RGB frames, ii) the full object might not be visible when the crop is taken based on hand poses only, and iii) hand crops have limitations in capturing global changes compared to the full frames. Hence, to capitalize on both the localized interaction information of hand crops and the global contextual information provided by full frames, our model combines them both. If a valid hand crop is found, we take the full and cropped RGB frames, pass them through the frame encoder, average their features, and re-normalize them to unit norm. This full vs. HOI crop ablation is shown in Tab. 13, in which combining both perform better than the alternatives.
@@ F Efficiency Comparison with Shift-GCN
While MS-G3D [43] and ISTA-Net [66] show state-of-the-art performance for action recognition with hand poses, they are not efficiency-focused. Our HandFormer outperforms them with significantly fewer FLOPs. However, HandFormer-B/6 prioritizes efficiency while slightly trading off accuracy. Therefore, we implement and test an efficiency-focused baseline, ShiftGCN [11], for verb recognition on Assembly101 [51] and compare it to HandFormer-B/6 in Tab. 14. While Shift-GCN relies on graph shift operations and pointwise convolutions for efficiency, our model identifies the crucial joints, i.e. fingertips and wrist joint, and processes only these joints to reduce FLOPs substantially. As evident from Tab. 14, our model outperforms Shift-GCN while incurring lower FLOPs.
@@ G.1 Pose-only Performance
Table 15: Action predictions by our model with and without sampling an RGB frame. Incorrect predictions are highlighted in red, while correct predictions are marked in green.

### v2

@@ 5.6 Ablation Studies
Keypoints. Not all joints of the hands are equally informative for understanding hand actions. For instance, fingertips exhibit greater mobility compared to the inner joints. Moreover, from an egocentric viewpoint, certain joints are more prone to self-occlusion than others. In Tab. 5, we present the impact of incorporating varying numbers of joints on verb recognition within the Assembly101 dataset [52]. For the case of 6 joints, we consider only the wrist joint along with the five fingertips. To expand to 11 joints, we incorporate all the joints along the index and the thumb, which are least affected by self-occlusion. We also show the effect of including the Global Wrist Token, which acts as a reference to the global motion of the hands while encoding micro-actions.
@@
Table 4: Egocentric action recognition in Assembly101 [52]. TSM features from [52] are used as RGB frame features.
Method Action Verb Object
TSM egocentric (fuse 4 views) 33.80 59.00 46.50
Egocentric (e3) + Pose 36.07 65.52 45.82
Egocentric (e4) + Pose 35.56 65.79 45.20
Egocentric (e3$+$e4) + Pose 38.05 66.32 47.86
Table 5: Keypoint ablation for verb recognition in Assembly101 [52]
 #Joints Global Verb
 Reference Accuracy (%)
21 64.17
21 64.90
11 64.77
6 63.70
@@
Micro-action length. As the resized input comprises a fixed number of pose frames, enlarging the window size for a micro-action decreases the number of micro-actions to aggregate, and vice-versa. In Tab. 6, we vary the micro-action length for verb recognition in the Assembly101 [52] using 6 joints per hand as input, i.e. fingertips and wrist joint. The input pose clip is temporally resized to $\mathcal{T}^{\prime}=120$ before breaking into micro-actions. Lengths 1 and 120 represent two extreme versions with frame-based and trajectory-based encoding, respectively, while the others conform to our micro-action-based formulation.
Table 6: Micro-action length ablation for verb recognition in Assembly101 [52]
#Frames 1 15 30 60 120
Verb  59.12 63.70  63.68  63.51  62.29
Accuracy(%)
Table 7: Ablating unimodal temporal aggregation for verb recognition in Assembly101 [52]
Temp. Agg. TCN LSTM Transformer
Verb Acc. (%)  62.95 63.34 63.70
@@
Temporal Aggregation. After extracting micro-action features, aggregation for action recognition can be done with any sequence model. In Tab. 7, we evaluate the effectiveness of different temporal aggregation methods for verb recognition in the Assembly101 dataset [52]. Here, we use 6 joints per hand as input, i.e., the fingertips and the wrist joint.
@@
Loss components. To assess the individual contributions of different components, we begin with a basic configuration. We then systematically introduce each element to understand its impact on the overall performance as observed in Tab. 8. Incorporating modality interaction between RGB and pose at the micro-action level through a multimodal tokenizer enhances action accuracy. The introduction of auxiliary losses also has a positive impact, resulting in an overall improvement of 2.08% for Assembly101 [52] and 7.44% for H2O [34].
Table 8: Ablating tokenization and different losses for action recognition
@@
 Multimodal  Feature  Verb & Object Action Accuracy (%) Tokenizer Ant. Loss Loss Assembly101 [52] H2O [34] 38.98 85.95 40.19 88.84 40.24 89.26 40.56 90.50 41.06 93.39
@@ 6 Conclusion
With the growing interest in AR/VR and wearables, hand pose estimation has rapidly advanced, and holds promise as a compact and domain-independent modality to complement the visual input. To address the underexplored domain of using 3D hand pose as a modality for hand-object interaction recognition, we introduce HandFormer, a novel multimodal transformer that leverages dense sequences of 3D hand poses with sparsely sampled RGB frames to achieve state-of-the-art action recognition performance. Our model also reduces computational requirements, offering immediate significance across various low-resource applications in mobile devices.
Limitations. Our method relies on the availability of hand poses, which, if extracted from the visual modality with pose estimation tools [27, 28], can encounter out-of-view scenarios and produce noisy poses. Our experiments reveal that these estimated poses can still achieve good accuracy, yet further research can be conducted to explicitly address this phenomenon. We also assume that a uniform sampling of RGB frames from each micro-action should provide good representations for understanding the semantic context. However, not all frames are equally important in understanding the action. In such a case, adaptive frame sampling methods can be employed, which we leave for future work.
Acknowledgements
This research is supported by A*STAR under its National Robotics Programme (NRP) (Award M23NBK0053).
@@ Acknowledgements
A Statistical Analysis: Full-body vs. Hand Skeletons
@@
This section serves as an extension of Sec. 3 to statistically analyze the differences between hand poses and full-body poses. Action recognition datasets [53, 41], which primarily focus on full-body actions, often include actions involving partial body movements. These actions exhibit limited global motion when viewed with respect to the entire skeleton, resulting in one or more relatively static joints. The change in the locations of moving joints with respect to such static joints can provide useful action cues. However, hand motions typically feature no such static reference points, as all the hand joints move together most of the time, making small changes in articulation to perform an action. To illustrate this difference, we randomly sample 1000 pose sequences from NTU RGB+D 120 [41] (full-body) and Assembly101 [52] (hands). We take the distances covered between two consecutive frames for each joint to form a distance array for the corresponding joint $j$ in a given pose sequence, which is determined by-
@@
For each of the selected 1000 sequences, we take the two temporal sequences $\{d_{j_{sta}}(t)\}_{t=1}^{T}$ and $\{d_{j_{dyn}}(t)\}_{t=1}^{T}$, normalize the distance values using the diameters of the corresponding skeletons, and plot the sequences separately in Fig. 7. As can be observed, compared to the distances covered by the most active joints in NTU RGB+D 120, the least active joints show significantly lower movement, effectively serving as the static reference points. On the other hand, the distinction in distance arrays between the most and the least active joints in Assembly101 is less pronounced. In addition, we calculated the Pearson correlation coefficient, denoted as $r$, between $\{d_{j_{sta}}(t)\}_{t=1}^{T}$ and $\{d_{j_{dyn}}(t)\}_{t=1}^{T}$ for all Assembly101 sequences, resulting in a value of 0.93. Conversely, for NTU RGB+D 120, the corresponding correlation coefficient is 0.33. This suggests strong coupling among hand joints during motion, and full-skeleton movement is more dominant in hand poses compared to full-body poses. Consequently, modeling dependencies between spatiotemporally distant joints is less effective for the highly dynamic hand motion (also discussed in Sec. 3). Therefore, by considering both long-term motion patterns and short-term articulation changes, our method facilitates efficient spatiotemporal factorization through micro-actions (Sec. 4.1). We also incorporate the full-skeletal motion from the entire action during micro-action encoding, using a global wrist token as a reference (Sec. 4.2).
@@
Fig. 8 illustrates sample heatmaps from NTU RGB+D 120 [41] and Assembly101 [52]. Keypoints in full-body human poses are often prominently situated, with minimal self-occlusion, and the subject is typically centered within the frame. As viewed in Fig. 8(a), reducing pose dimensions to 2D does not significantly compromise detail; rather, it enhances input reliability by simplifying the pose estimation problem. However, this advantage diminishes when applied to hand poses. Hand poses present unique challenges, such as frequent self-occlusion and closer proximity of the keypoints, which are exacerbated by reducing the dimension to 2D. To empirically analyze this phenomenon, we evaluate HandFormer-B/6 with 2D and 3D poses for recognizing verbs in Assembly101 [52] and report in Tab. 9. This analysis reveals about $5$% difference in favor of the 3D input. Furthermore, PoseConv3D [20] introduces a CNN-based approach with 2D keypoints, which directly utilizes heatmaps from the pose estimator or generates Gaussian heatmaps from the 2D coordinates. However, feeding heatmaps can diminish the clarity of keypoints to the model, particularly when they are in close proximity, as is often the case with hand poses. Hence, PoseConv3D [20] performs poorly in recognizing hand actions, as evident in Tab. 9.
@@
C Alternatives for Frame Encoder
We utilize pre-trained ViT-g/14 and ViT-L/14 models from DINOv2 [47] followed by a linear layer without any fine-tuning as the frame encoder $F$ for Assembly101 [52] and H2O [34], respectively. Except for our experiments on Assembly101 [52] with monochrome egocentric videos (Sec. 5.5), all our multimodal results are obtained using DINOv2 features. This approach enables us to assess the effectiveness of image-based foundation models in videos, leveraging all-purpose features from RGB frames and achieving strong cross-view generalization performance (Sec. 5.4). Although this is our default choice offering easy adaptation to new datasets with faster training, it can be considered compute-heavy when deployed in a low-resource setting due to the large ViT backbones. For better efficiency during inference, we propose a pretraining scheme that allows us to use a ResNet50 [29] that replaces the ViT without compromising accuracies, as shown in Tab. 10. Specifically, we first train a TSM [40] model with a ResNet50 [29] backbone for action recognition, utilizing all action clips and then dropping the classification layer. This ResNet50 becomes the frozen image encoder in our proposed architecture, replacing ViT. During the training and inference of HandFormer, the TSM backbone operates as a true image model (ResNet50), as we employ it on individual frames without any channel shifting. The TSM features provided in the Assembly101 [52] are generated in this way, and we utilize them in our egocentric action recognition experiments (Sec. 5.5).
@@ C Alternatives for Frame Encoder
D Maintaining High Temporal Resolution at Low Cost
Our method is designed to perform action recognition efficiently in hand-object interaction videos. Obtaining efficiency in such a setup is challenging as we need to maintain a high temporal resolution to understand fine-grained hand movements that constitute the actions. Therefore, we propose HandFormer using densely sampled pose frames and sparse RGB frames. In this section, we quantify the efficiency of this method compared to an alternative video model. As mentioned, understanding fine-grained hand motion demands a high temporal resolution to differentiate verb classes. For instance, relying on sparsely sampled frames may make actions like “screwing" and “unscrewing" indistinguishable. However, adopting a high temporal resolution with video models operating on RGB frames is challenging, primarily due to (i) the excessive computation associated with performing spatiotemporal operations on numerous frames, and (ii) the need to address redundancy in RGB frames to extract meaningful information.
@@ E.2 Hand-Object Interaction Crop
In hand-object interaction (HOI) videos, the region of interest typically centers around the hands, capturing crucial information about the interacting object and the type of interaction. Leveraging 3D hand poses obtained through a readily available pose estimator [27], we project these poses onto RGB frames, extract the enclosing rectangle of the projected 2D pose, and expand it by 25% to define the ROI crop. However, relying solely on the cropped region can occasionally mislead the model for three potential reasons: i) failure of the pose estimator on certain frames, leading to the absence of useful features from the RGB frames, ii) the full object might not be visible when the crop is taken based on hand poses only, and iii) hand crops have limitations in capturing global changes compared to the full frames. Hence, to capitalize on both the localized interaction information of hand crops and the global contextual information provided by full frames, our model combines them both. If a valid hand crop is found, we take the full and cropped RGB frames, pass them through the frame encoder, average their features, and re-normalize them to unit norm. This full vs. HOI crop ablation is shown in Tab. 13, in which combining both performs better than the alternatives.
@@ F Efficiency Comparison with Shift-GCN
While MS-G3D [44] and ISTA-Net [66] show state-of-the-art performance for action recognition with hand poses, they are not efficiency-focused. Our HandFormer outperforms them with significantly fewer FLOPs. However, HandFormer-B/6 prioritizes efficiency while slightly trading off accuracy. Therefore, we implement and test an efficiency-focused baseline, ShiftGCN [11], for verb recognition on Assembly101 [52] and compare it to HandFormer-B/6 in Tab. 14. While Shift-GCN relies on graph shift operations and pointwise convolutions for efficiency, our model identifies the crucial joints, i.e., the fingertips and the wrist joint, and processes only these joints to reduce FLOPs substantially. As evident from Tab. 14, our model outperforms Shift-GCN while incurring lower FLOPs.
@@ G.1 Pose-only Performance
Table 15: Action classification by our model with and without sampling an RGB frame. Incorrect predictions are highlighted in red, while correct predictions are marked in green.